Grade 12th passAlgebra

To find the square root of the product of 998×999×1000×1001×1002

Sambit Dash
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
the square rootCheck this first: and get an idea how to solve:
of √(999x1000x1001x1002+1)
 
 
 
Let y = n2+3n
 
Then, the given expression is y(y+2)+1 = y2+2y+1 = (y+1)2 = (n2+3n+1)2.
 
Hence, the square root is n2+3n+1 . Substitue n = 999, and you get the square root as 1000999